For a particular device, in the house, it has alternating current (AC) and voltage of 20.0 A and 280 V respectively. What are their amplitude values?




1
Expert's answer
2021-07-05T08:44:53-0400

Since we know that "V_{rms}=\frac{1}{\sqrt{2}}V_{max}" and "I_{rms}=\frac{1}{\sqrt{2}}I_{max}" , we find the amplitud values for both magnitudes as:


"V_{max}=\sqrt{2}V_{rms}=\sqrt{2}(280\,V)=395.98\,V
\\ I_{max}=\sqrt{2}I_{rms}=\sqrt{2}(20.0\,A)=28.28\,A"


In conclusion, the amplitude of the voltage is 395.98 V and the amplitude for the current is 28.28 A.


Reference:

  • Young, H. D., & Freedman, R. A. (2015). University Physics with Modern Physics and Mastering Physics (p. 1632). Academic Imports Sweden AB.